OpenStudy (anonymous):

The distance between two locations, A and B, is calculated using a third location C at a distance of 15 miles from location B. If B = 105° and C = 20°, what is the distance, to the nearest tenth of a mile, between locations A and B? A 42.4 miles B 35.9 miles C 6.3 miles D 5.3 miles

OpenStudy (anonymous):

So I made a triangle:  To solve for X, we can fall back on our useful little term, cosine. Here. Lemme show ya how it works. Cosine (of some given angle) = Adjacent Side / Hypotenuse. We know this. Soooooo, let's fill in teh blanks. Cos (20) = (15) / (x). This is the same thing as saying: (x) = ( (15) / (cos)(20) ). When you solve dat out, you get: x = 15.96. So now we've solved for X. Pretty easy to solve for Y now, right? Right. Just use the Pythagorean Theorem. Fill in the X-side on my little triangle diagram to help you visualize what's happening a bit better. The longest side of our constructed triangle is equal to 15.96, so: (15.96)^2 = (y)^2 + (15)^2 We'll solve that out. (15.96)^2 = (y)^2 + (15)^2 ...Becomes: (254.81) = (y)^2 + (225) ...Becomes: (y)^2 = 29.81 ...Becomes: (y) = 5.46. This value here represents the distance from Point A to Point B. The closest choice to this answer is: D.